OpenStudy (anonymous):

help plz for medal and a fan!!!!! Twelve and seventeen-eighteenths minus 4 and two-thirds is equal to what number? 8 and five-sixths 8 and five-thirteenths 8 and five-eighteenths 3 and one-sixteenth

OpenStudy (anonymous):

Do you know the first step in solving this

OpenStudy (anonymous):

Or do you have no clue...I will be happy to explain it

OpenStudy (anonymous):

@zaibali.qasmi

OpenStudy (anonymous):

help @zaibali.qasmi

OpenStudy (anonymous):

just a moment...

OpenStudy (anonymous):

\[=12\times \frac{ 17 }{ 18 }-4\times \frac{ 2 }{ 3 }=\frac{ 233 }{ 18 }-\frac{ 14 }{ 3 }\] taking LCM \[=\frac{ 233-84 }{ 18 }=\frac{ 149 }{ 18 } =8.277= 8 and five-eighteenths

OpenStudy (anonymous):

8 and 5 18 is the answer?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Wcm///